# Simultaneous Inference in Multiple Matrix-Variate Graphs for High-Dimensional Neural Recordings

### v1

@@ Abstract
As large-scale neural recordings become common, many neuroscientific investigations are focused on identifying functional connectivity from spatio-temporal measurements in two or more brain areas across multiple sessions. Spatial-temporal data in neural recordings can be represented as matrix-variate data, with time as the first dimension and space as the second. In this paper, we exploit the multiple matrix-variate Gaussian Graphical model to encode the common underlying spatial functional connectivity across multiple sessions of neural recordings. By effectively integrating information across multiple graphs, we develop a novel inferential framework that allows simultaneous testing to detect meaningful connectivity for a target edge subset of arbitrary size. Our test statistics are based on a group penalized regression approach and a high-dimensional Gaussian approximation technique. The validity of simultaneous testing is demonstrated theoretically under mild assumptions on sample size and non-stationary autoregressive temporal dependence. Our test is nearly optimal in achieving the testable region boundary. Additionally, our method involves only convex optimization and parametric bootstrap, making it computationally attractive. We demonstrate the efficacy of the new method through both simulations and an experimental study involving multiple local field potential (LFP) recordings in the Prefrontal Cortex (PFC) and visual area V4 during a memory-guided saccade task.
@@ 1 Introduction
We address the problem of discovering associations between simultaneously recorded spatiotemporal datasets. A motivating example is neural recordings captured by two Utah arrays in a monkey’s brain. Each Utah array comprises a $10\times 10$ square grid of electrodes on a $4.2\mathrm{mm}\times 4.2\mathrm{mm}$ base (Jones et al.,, 1992), and each electrode records pooled electrical signals or local field potentials from nearby neurons (Buzsáki et al.,, 2012). Khanna et al., (2020) recorded neural activity in both the visual (V4) and prefrontal (PFC) cortices during a spatial working memory task, producing 100 time series from each brain region. The coordinated activity between these two sets of time series can reveal the relationship between the two spatially-distant regions, or the functional connection.
When each dataset consists of a single time series, the coordinated activity between two datasets can be quantified using cross-correlation, calculated based on the replication structure within the datasets (Gubner,, 2006). In neural experiments, each replication is referred to as an experimental trial, and a group of trials forms a session. While estimation precision increases with more trials, the duration of a single session is often constrained by physical, economic, and ethical factors. These limitations require pooling observations across multiple sessions, but the primary statistical challenge is accounting for the heterogeneity between trials from different sessions. Another challenge is the summarization of cross-correlations between all possible pairs of time series in the two spatiotemporal datasets. For instance, in a dataset with $100$ time series from each brain region, there are $10,000$ cross-regional pairs. To address this, we propose a novel statistical inference approach for the V4-PFC connection by assessing the global significance of these cross-regional interactions. We achieve this by employing a multiple matrix-variate Gaussian graphical model, where the spatial and temporal partial correlation parameters represent the association structure within the spatiotemporal datasets for each session.
Building on the matrix-variate Gaussian graphical model outlined above, we propose a three-step procedure for estimating and inferring the spatial partial correlations. First, we estimate the spatial partial correlation parameters across multiple sessions using a node-wise regression approach (Ren et al.,, 2019). Next, we estimate the temporal covariance parameters, assuming a mild bandable structure on their Cholesky factors, which will be used as nuisance parameters in the final inference step. Finally, leveraging the results from the first two steps, we construct a novel test statistic to evaluate the significance of individual edges or groups of edges. This inference framework draws inspiration from the linear functional-based tests of Ren et al., (2019), the debiased inference techniques for matrix-variate graphs from Chen and Liu, (2018), and the high-dimensional Gaussian approximation methods by Chernozhukov et al., 2023b .
In addition, our work provides several methodological advancements and theoretical contributions. First, we develop a method for leveraging multiple session data collectively based on group lasso, enabling the discovery of a common spatial correlation graph across sessions with improved sensitivity. We demonstrate the improved inferential performance in theoretical analyses (Section 4) and simulation studies (Section 5). Second, while group lasso is a well-known method, its application and consistency in the context of auto-correlated samples have not been studied theoretically. In Section 4, we present a self-contained analysis, deriving rates of convergence for both estimation and prediction. This contribution may have broader implications for high-dimensional statistics involving correlated data. Finally, we provide theoretical guarantees for our proposed inference procedures, employing a Gaussian approximation bootstrap method based on Chernozhukov et al., 2023b . Notably, this is the first formulation of a global significance test for a group of edges in partial-correlation graphs, distinguishing our approach from existing FDR control methods.
1.1 Notations
@@
For a vector $\mathbf{x}$, let $\lVert\mathbf{x}\rVert_{p}$ denote the $\ell_{p}$-norm. For a real matrix $X$, we use $X_{i,\cdot}$ to denote the $i$-th row of $X$ and $X_{\cdot,j}$ to denote the $j$-th column of $X$. Let $\lVert X\rVert_{p}$ denote the matrix $p$-norm, where $\lVert X\rVert_{2}$ is also referred to as the spectral or operator norm. We denote the Frobenius norm of $X$ by $\lVert X\rVert_{F}$ and the entry-wise supremum norm by $\lVert X\rVert_{\infty}$. For any set $A$, its cardinality is denoted by $\lvert A\rvert$.
@@ 2 Matrix-variate Gaussian Graphical Model
Let $X\in\mathbb{R}^{p\times q}$ be the observed spatiotemporal data in a single trial. At a given time point $t$, $X_{t,\cdot}\in\mathbb{R}^{q}$ is vector-valued, with each entry of the vector corresponding to a specific spatial point. Vector-valued graphical models have been extensively employed to explore conditional dependence relationships in high-dimensional data, including neural recordings (Fornito et al.,, 2013; Vinci et al.,, 2018). In these models, each node in an undirected graph represents a vector entry, and two nodes are connected by an edge if and only if their corresponding entries exhibit conditional dependence given all other entries. When the random vector follows a multivariate Gaussian distribution, the partial correlations, or equivalently elements of the precision matrix, encode conditional dependencies between entries (Meinshausen and Bühlmann,, 2006). Consequently, this type of graph is also referred to as a partial-correlation graph.
Since spatiotemporal data involve both time ($t=1,\dots,p$) and space ($i=1,\dots,q$), we consider a matrix-variate extension. A matrix-variate Gaussian graphical model (Dawid,, 1981) is characterized by two covariance parameters: temporal (row) covariance matrix $\Sigma^{(\mathcal{T})}\in\mathbb{R}^{p\times p}$ and spatial (column) covariance matrix $\Sigma^{(\mathcal{S})}\in\mathbb{R}^{q\times q}$. The observed data $X$ are said to follow the model with mean $\mu$, denoted by $\operatorname{MN}(\mu,\Sigma^{(\mathcal{T})},\Sigma^{(\mathcal{S})})$, if and only if the vectorized matrix $\mathrm{vec}(X)$ is vector-variate Gaussian distributed with mean $\mathrm{vec}(\mu)$ and variance-covariance matrix $\Sigma^{(\mathcal{S})}\otimes\Sigma^{(\mathcal{T})}$, where $\otimes$ denotes the Kronecker product. Due to the identifiability issue, hereafter we always assume that $\operatorname{tr}(\Sigma^{(\mathcal{T})})=p$ on the temporal (row) covariance. By imposing this Kronecker product structure, we reduce the number of parameters on the covariance from the order of $p^{2}q^{2}$ to $p^{2}+q^{2}$, while achieving an interpretable covariance structure for both time and space. Indeed, this structure implies that the covariance between $X_{ti}$ and $X_{sj}$ is $\Sigma^{(\mathcal{T})}_{ts}\cdot\Sigma^{(\mathcal{S})}_{ij}$. If we denote the temporal (row) and spatial (column) precision matrices by $\Omega^{(\mathcal{T})}=(\Sigma^{(\mathcal{T})})^{-1}$ and $\Omega^{(\mathcal{S})}=(\Sigma^{(\mathcal{S})})^{-1}$, respectively, the inverse of $\Sigma^{(\mathcal{S})}\otimes\Sigma^{(\mathcal{T})}$ has a simple analytic form $(\Sigma^{(\mathcal{S})}\otimes\Sigma^{(\mathcal{T})})^{-1}=\Omega^{(\mathcal{S})}\otimes\Omega^{(\mathcal{T})}$. Due to the relationship between partial correlation and precision matrix, the partial correlation between space points $i$ and $j$ is given by $\rho^{(\mathcal{S})}_{ij}=-{\Omega^{(\mathcal{S})}_{ij}}/{\sqrt{\Omega^{(\mathcal{S})}_{ii}\Omega^{(\mathcal{S})}_{jj}}},$ invariant across time points (see Section 3.1).
@@ 3 Simultaneous Inference Framework
For each session $l=1,\dots,m$, we observe spatiotemporal measurements in $p\times q$ matrix-variate sample over We assume each sample $X^{(k,l)}$ at trial $k$ and session $l$ follows matrix-variate Gaussian distribution $\operatorname{MN}(0,\Sigma^{(\mathcal{T},l)},\Sigma^{(\mathcal{S},l)})$, where $\Sigma^{(\mathcal{T},l)}\in\mathbb{R}^{p\times p}$ and $\Sigma^{(\mathcal{S},l)}\in\mathbb{R}^{q\times q}$ are the session-specific temporal and spatial covariance matrices, respectively. We denote the temporal and spatial precision matrices by $\Omega^{(\mathcal{T},l)}:=(\Sigma^{(\mathcal{T},l)})^{-1}$ and $\Omega^{(\mathcal{S},l)}:=(\Sigma^{(\mathcal{S},l)})^{-1}$, respectively.
@@
1:Input: Multi-session data $\mathcal{D}$, edge set $\mathcal{S}$, test level $\alpha$
@@ 3.1 Estimation of Spatial Covariance Matrix
We begin by considering a single session $l$ and time point $t$ to motivate our node-wise regression approach (Meinshausen and Bühlmann,, 2006; Liu,, 2013). For the random variable $X^{(k,l)}_{ti}$, which represents electrode $i$ at time $t$, its conditional distribution given the remaining variables in $X^{(k,l)}_{t,\cdot}$ follows a normal distribution and can be modeled as a linear regression:
@@
where the regression coefficients are related to the spatial precision matrix via $\beta^{(\mathcal{S},l)}_{ji}=-\frac{\Omega^{(\mathcal{S},l)}_{ji}}{\Omega^{(\mathcal{S},l)}_{ii}}\mathbb{I}(i\neq j),$ and $\mathbb{E}[\epsilon^{(\mathcal{S},k,l)}_{ti}]=0$. Due to the Kronecker product structure in the model, the regression coefficients are time-independent. Consequently, sparsity in the spatial precision matrix $\Omega^{(\mathcal{S},l)}$ implies corresponding sparsity in the regression coefficients $\beta^{(\mathcal{S},l)}$. A similar relationship holds between the spatial precision matrix and the covariance of the node-wise regression residuals. Specifically, for the residuals $\epsilon^{(\mathcal{S},k,l)}_{t,\cdot}=(\epsilon^{(\mathcal{S},k,l)}_{t1},\dots,\epsilon^{(\mathcal{S},k,l)}_{tq})^{\top}$, the covariance is given by $\operatorname{Cov}[\epsilon^{(\mathcal{S},k,l)}_{t,\cdot}]=\Sigma^{(\mathcal{T},l)}_{tt}\cdot\Phi^{(\mathcal{S},l)},$ where $\Phi^{(\mathcal{S},l)}$ is a time-independent matrix with elements defined as $\Phi^{(\mathcal{S},l)}_{ij}=\frac{\Omega^{(\mathcal{S},l)}_{ij}}{\Omega^{(\mathcal{S},l)}_{ii}\Omega^{(\mathcal{S},l)}_{jj}}$. Our target, the spatial partial correlation, is expressed as $\rho^{(\mathcal{S},l)}_{ij}=-\frac{\Omega^{(\mathcal{S},l)}_{ij}}{\sqrt{\Omega^{(\mathcal{S},l)}_{ii}\Omega^{(\mathcal{S},l)}_{jj}}}=-\frac{\Phi^{(\mathcal{S},l)}_{ij}}{\sqrt{\Phi^{(\mathcal{S},l)}_{ii}\Phi^{(\mathcal{S},l)}_{jj}}},$ and testing $\Phi^{(\mathcal{S},l)}_{ij}=0$ is equivalent to testing whether the partial correlation between $i$ and $j$ is zero. In Eq. 1, we treat each row of $X^{(k,l)}$ as a $q$-dimensional sample, giving us $p$ correlated vector-valued samples for a sparse linear regression model, where the covariance among these “row samples” is characterized by $\Sigma^{(\mathcal{T},l)}$. Collecting all trials $k=1,\dots,n_{l}$, we estimate $\Phi^{(\mathcal{S},l)}$ using $\Phi^{(\mathcal{S},l)}=\mathbb{E}\left[\frac{1}{n_{l}p}\sum_{k=1}^{n_{l}}\sum_{t=1}^{p}\epsilon^{(\mathcal{S},k,l)}_{t,\cdot}\epsilon^{(\mathcal{S},k,l)\top}_{t,\cdot}\right],$ under the identifiability constraint $\operatorname{tr}(\Sigma^{(\mathcal{T},l)})=p$, where we effectively use $n_{l}p$ correlated samples in total.
@@
Having discussed the model for individual sessions, we now consider all $m$ sessions jointly to improve the estimation accuracy of each $\rho^{(\mathcal{S},l)}_{ij}$. Under assumption that the spatial precision matrices $\Omega^{(\mathcal{S},l)}_{ij}$ share the same sparsity pattern across sessions, the corresponding regression coefficients $\beta^{(\mathcal{S},l)}_{\cdot,i}$ also exhibit a common support across the $m$ sessions. To harness this group sparsity assumption, we treat the coefficients $\beta^{(\mathcal{S},1)}_{ij},\dots,\beta^{(\mathcal{S},m)}_{ij}$ as a group of parameters for each pair $(i,j)$ and apply group lasso (Yuan and Lin,, 2006) to the stacked $m$ linear models, thereby improving estimation by borrowing strength across sessions:
$\begin{split}&\{\widehat{\beta}^{(\mathcal{S},l)}_{\cdot,i}\}_{l=1,\dots,m}:=\\ &\operatorname*{\arg\!\min}_{\{b^{(l)}\}_{l=1,\dots,m}}\left\{\frac{1}{2n_{0}p}\sum_{l=1}^{m}\sum_{k=1}^{n_{l}}\lVert X^{(k,l)}_{\cdot,i}-X^{(k,l)}b^{(l)}\rVert_{2}^{2}+\gamma_{i}\sum_{j:j\neq i}\sqrt{\sum_{l=1}^{m}\frac{\lVert\sum_{k}X^{(k,l)}_{\cdot,j}\rVert_{2}^{2}}{n_{l}p}b^{(l)2}_{j}}\right\},\\ \end{split}$ (2)
@@
Once the regression coefficients are estimated, the fitted residuals are computed as $\widehat{\epsilon}^{(\mathcal{S},k,l)}_{ti}=X^{(k,l)}_{ti}-X^{(k,l)}_{t,\cdot}\widehat{\beta}^{(\mathcal{S},l)}_{\cdot,i}$. Although the empirical covariance matrix of the fitted residuals provides a straightforward estimate of $\Phi_{ij}^{(\mathcal{S},l)}$, this estimate is biased due to the lasso-type penalty, leading to a larger error rate than the expected $1/\sqrt{n_{l}p}$ rate (Liu,, 2013). To address this bias, we introduce a bias-correction term in the covariance estimate:
@@
Using this bias-corrected estimate, we then estimate the spatial precision matrix $\Omega^{(\mathcal{S},l)}$ and the partial correlation matrix $\rho^{(\mathcal{S},l)}$ as follows:
@@
While our partial correlation estimator is similar to the form proposed by Chen and Liu, (2018) for a single matrix-variate Gaussian graphical model, our approach leverages information from multiple sessions using the group lasso estimate, leading to a faster convergence rate, as demonstrated in Theorem 4.1.

### v2

@@ Abstract
We study simultaneous inference for multiple matrix-variate Gaussian graphical models in high-dimensional settings. Such models arise when spatiotemporal data are collected across multiple sample groups or experimental sessions, where each group is characterized by its own graphical structure but shares common sparsity patterns. A central challenge is to conduct valid inference on collections of graph edges while efficiently borrowing strength across groups under both high-dimensionality and temporal dependence. We propose a unified framework that combines joint estimation via group penalized regression with a high-dimensional Gaussian approximation bootstrap to enable global testing of edge subsets of arbitrary size. The proposed procedure accommodates temporally dependent observations and avoids naive pooling across heterogeneous groups. We establish theoretical guarantees for the validity of the simultaneous tests under mild conditions on sample size, dimensionality, and non-stationary autoregressive temporal dependence, and show that the resulting tests are nearly optimal in terms of the testable region boundary. The method relies only on convex optimization and parametric bootstrap, making it computationally tractable. Simulation studies and a neural recording example illustrate the efficacy of the proposed approach.
@@ 1 Introduction
Graphical models provide a powerful framework for characterizing dependency structures in high-dimensional data. Their matrix-variate extensions have been widely used for analyzing high-dimensional spatiotemporal data, with prominent applications including functional connectivity networks among brain regions in neuroscience (Kass et al., 2023; Bong et al., 2021) and financial stock price networks in econometrics (Chang et al., 2018). A central challenge in these settings arises from the high dimensionality of the target graph, where the prohibitive number of candidate edges leads to a severe multiple testing burden. Addressing this challenge requires not only new inferential tools but also careful and efficient use of available data. In many applications, it is necessary to combine data from multiple sample groups collected under heterogeneous conditions, such as different subjects or experimental sessions, in order to overcome the limited sample size available within each group. However, naive pooling across groups can introduce spurious dependencies due to heterogeneity, while analyzing each group separately often results in substantial loss of statistical power. In this paper, we develop a method for jointly estimating multiple matrix-variate graphical models by leveraging shared sparsity structures across sample groups, together with a unified framework for high-dimensional inference on graph edges.
The literature on inference for graphical models is extensive (Friedman et al., 2008; Allen and Tibshirani, 2010; Jankova and van de Geer, 2015; Chen and Liu, 2018; Ye et al., 2021; Liu, 2013; Ren et al., 2015), but relatively few frameworks address matrix-variate Gaussian graphical models (MGGMs; see Section 2). Existing work includes (Leng and Tang, 2012; Yin and Li, 2012; Zhou, 2014; Chen and Liu, 2018; Ye et al., 2021). For example, Chen and Liu (2018) developed a multiple testing framework for support recovery with asymptotic guarantees on normality and false discovery rate control, while Ye et al. (2021) proposed paired tests for detecting changes in matrix graph across correlated samples. Both approaches rely on multiple testing procedures with false discovery rate control. In contrast, Chang et al. (2018) derived a single global test statistic for assessing subgroups of edges simultaneously in vector-variate graphical models under temporal dependence.
Inference procedures for multiple matrix-variate graphs remain largely unexplored. In the vector-variate setting, several approaches have been proposed for joint estimation of multiple graphs, including optimization-based methods (Danaher et al., 2014; Cai et al., 2016a; Lee and Liu, 2015) and Bayesian formulations with structured priors (Peterson et al., 2015). The work most closely related to our setting is Zhu and Li (2018), who studied estimation of multiple MGGMs via nonconvex optimization with sparsity and group Lasso penalties. While their estimator exhibits some desirable properties, its convergence relies on strong assumptions and formal inference procedures are not developed.
We develop and study a multiple matrix-variate graph-based method for discovering common sparsity structures with improved sensitivity. Our approach uses group Lasso to borrow strength across sample groups, yielding more accurate partial-correlation estimates than the naive strategy of fitting models separately to each group; see Section 3 for details. We establish improved convergence rates through theoretical analysis and simulation studies in Sections 4 and 5. In particular, we provide a self-contained analysis of the convergence rates and prediction error of the group Lasso estimator under temporal dependence in Section 4.1. Although group Lasso itself dates to 2006 (Yuan and Lin, 2006), its theoretical properties under temporal correlation have not previously been studied, and may be of independent interest for high-dimensional statistics with dependent data. In Section 4.2, we develop theoretical guarantees for the proposed inference procedures using a Gaussian approximation bootstrap method based on Chernozhukov et al. (2013). This represents the first attempt to formulate a global significance test for edge groups of arbitrary size in matrix-variate graphs beyond false discovery rate based methods. Finally, we apply our method to microelectrode recordings from two brain regions, V4 and PFC, and identify distinct cross-area subgroups with differential dynamics across cognitive stages (Section 5.2). Our results reveal increased V4–PFC connectivity during visual memory retention, supporting their joint role in working memory.
@@
Notations: For a vector $\mathbf{x}$, let $\lVert\mathbf{x}\rVert_{p}$ denote the $\ell_{p}$-norm. For a real matrix $X$, we use $X_{i,\cdot}$ to denote the $i$-th row of $X$ and $X_{\cdot,j}$ to denote the $j$-th column of $X$. Let $\lVert X\rVert_{p}$ denote the matrix $p$-norm, where $\lVert X\rVert_{2}$ is also referred to as the spectral or operator norm. We denote the Frobenius norm of $X$ by $\lVert X\rVert_{F}$ and the entry-wise supremum norm by $\lVert X\rVert_{\infty}$. For a set $A$, its cardinality is denoted by $\lvert A\rvert$. We denote the zero vector, whose elements are all zero, by $\mathbf{0}$. For a positive integer $d$, we use $[d]$ to denote the set $\{1,2,...,d\}$.
@@ 2 Matrix-variate Gaussian Graphical Model
Let $X\in\mathbb{R}^{p\times q}$ denote an observed sample of spatiotemporal data. At a given time point $t\in[p]$, the row vector $X_{t,\cdot}\in\mathbb{R}^{q}$ represents measurements across $q$ spatial locations. A matrix-variate Gaussian graphical model (MGGM, Dawid, 1981) is specified by two covariance matrices: the temporal (row) covariance $\Sigma^{(\mathcal{T})}\in\mathbb{R}^{p\times p}$ and the spatial (column) covariance $\Sigma^{(\mathcal{S})}\in\mathbb{R}^{q\times q}$. We say that $X$ follows this model with mean $\mu$, denoted by $X\sim\mathrm{MN}(\mu,\Sigma^{(\mathcal{T})},\Sigma^{(\mathcal{S})}),$ if and only if the vectorized data $\mathrm{vec}(X)$ follow a multivariate normal distribution with mean $\mathrm{vec}(\mu)$ and covariance matrix $\Sigma^{(\mathcal{S})}\otimes\Sigma^{(\mathcal{T})}$, where $\otimes$ denotes the Kronecker product.
Because this parametrization is not identifiable up to a scaling factor, we assume without loss of generality that $\operatorname{tr}(\Sigma^{(\mathcal{T})})=p$. Imposing the Kronecker product structure reduces the number of free covariance parameters from $pq(pq+1)/2$ to $p(p+1)/2+q(q+1)/2$. Under this model, the covariance between entries $X_{ti}$ and $X_{sj}$ is given by $\operatorname{Cov}(X_{ti},X_{sj})=\Sigma^{(\mathcal{T})}_{ts}\,\Sigma^{(\mathcal{S})}_{ij}$. If we denote the corresponding precision matrices by $\Omega^{(\mathcal{T})}:=(\Sigma^{(\mathcal{T})})^{-1}$ and $\Omega^{(\mathcal{S})}:=(\Sigma^{(\mathcal{S})})^{-1}$, then the inverse covariance of $\mathrm{vec}(X)$ admits a simple analytic form: $(\Sigma^{(\mathcal{S})}\otimes\Sigma^{(\mathcal{T})})^{-1}=\Omega^{(\mathcal{S})}\otimes\Omega^{(\mathcal{T})}.$ In MGGM, interest primarily lies in the conditional independence graph structure among $q$ spatial locations, where each node represents a spatial location and the absence of an edge indicates that the corresponding nodes are conditionally independent given all remaining nodes. For Gaussian distributions, this graph structure is encoded by the precision matrix, or equivalently, by partial correlations. Thus, the target parameters in this work are the spatial partial correlations, which can be recovered from the spatial precision matrix: the partial correlation between spatial points $i$ and $j$ is $\rho^{(\mathcal{S})}_{ij}=-\frac{\Omega^{(\mathcal{S})}_{ij}}{\sqrt{\Omega^{(\mathcal{S})}_{ii}\,\Omega^{(\mathcal{S})}_{jj}}}$, which remains invariant across all time points (see Section 3.1).
@@ 3 Simultaneous Inference Framework
For each sample group $l\in[m]$, we observe $n_{l}$ i.i.d. spatiotemporal measurements in $p\times q$ matrix-variate samples. At group $l$, we assume for $k\in[n_{l}]$ that the $k$-th sample $X^{(k,l)}$ follows matrix-variate Gaussian distribution $\operatorname{MN}(0,\Sigma^{(\mathcal{T},l)},\Sigma^{(\mathcal{S},l)})$, where $\Sigma^{(\mathcal{T},l)}\in\mathbb{R}^{p\times p}$ and $\Sigma^{(\mathcal{S},l)}\in\mathbb{R}^{q\times q}$ are the group-specific temporal and spatial covariance matrices, respectively. We denote the temporal and spatial precision matrices by $\Omega^{(\mathcal{T},l)}:=(\Sigma^{(\mathcal{T},l)})^{-1}$ and $\Omega^{(\mathcal{S},l)}:=(\Sigma^{(\mathcal{S},l)})^{-1}$, respectively.
@@
1:Input: Multi-group data $\mathcal{D}$, edge set $\mathcal{S}$, test level $\alpha$
@@ 3.1 Estimation of Spatial Covariance Matrix
Under the matrix-variate Gaussian graphical model, for a given spatial location $i\in[q]$, at each time point $t\in[p]$, the conditional distribution of the random variable $X^{(k,l)}_{ti}$ given the remaining variables in $X^{(k,l)}_{t,\cdot}$ follows a normal distribution and can be modeled as a linear regression:
@@
where the regression coefficients are related to the spatial precision matrix via $\beta^{(\mathcal{S},l)}_{ji}=-\frac{\Omega^{(\mathcal{S},l)}_{ji}}{\Omega^{(\mathcal{S},l)}_{ii}}\mathbb{I}(i\neq j),$ and $\mathbb{E}[\epsilon^{(\mathcal{S},k,l)}_{ti}]=0$. Due to the Kronecker product structure in the model, the regression coefficients are time-independent. It is common in many applications of graphical models to assume that the graph exhibits a certain sparsity structure. Consequently, sparsity in the spatial precision matrix $\Omega^{(\mathcal{S},l)}$ implies corresponding sparsity in the regression coefficients $\beta^{(\mathcal{S},l)}$. A similar sparsity appears in the covariance of the node-wise regression residuals. Specifically, for the residuals $\epsilon^{(\mathcal{S},k,l)}_{t,\cdot}:=(\epsilon^{(\mathcal{S},k,l)}_{t1},\dots,\epsilon^{(\mathcal{S},k,l)}_{tq})^{\top}$, the covariance is given by $\operatorname{Cov}[\epsilon^{(\mathcal{S},k,l)}_{t,\cdot}]=\Sigma^{(\mathcal{T},l)}_{tt}\cdot\Phi^{(\mathcal{S},l)},$ where $\Phi^{(\mathcal{S},l)}_{ij}:=\frac{\Omega^{(\mathcal{S},l)}_{ij}}{\Omega^{(\mathcal{S},l)}_{ii}\Omega^{(\mathcal{S},l)}_{jj}}$. Our target, the spatial partial correlation, is expressed as $\rho^{(\mathcal{S},l)}_{ij}=-\frac{\Omega^{(\mathcal{S},l)}_{ij}}{\sqrt{\Omega^{(\mathcal{S},l)}_{ii}\Omega^{(\mathcal{S},l)}_{jj}}}=-\frac{\Phi^{(\mathcal{S},l)}_{ij}}{\sqrt{\Phi^{(\mathcal{S},l)}_{ii}\Phi^{(\mathcal{S},l)}_{jj}}}$, and testing $\Phi^{(\mathcal{S},l)}_{ij}=0$ is equivalent to testing $\rho^{(\mathcal{S},l)}_{ij}=0$. Because $\Phi^{(\mathcal{S},l)}$ is invariant across $t$ and $k$, $\Phi^{(\mathcal{S},l)}=\mathbb{E}\left[\frac{1}{n_{l}p}\sum_{k=1}^{n_{l}}\sum_{t=1}^{p}\epsilon^{(\mathcal{S},k,l)}_{t,\cdot}\epsilon^{(\mathcal{S},k,l)\top}_{t,\cdot}\right$], noting that $\operatorname{tr}(\Sigma^{(\mathcal{T},l)})=p$ under the identifiability constraint. It is worth noting that the residuals corresponding to different time points $t\in[p]$ exhibit temporal dependence.
@@
We fit all $m$ sample groups jointly to improve the estimation accuracy of each $\rho^{(\mathcal{S},l)}_{ij}$. Under a mild assumption that the spatial precision matrices $\Omega^{(\mathcal{S},l)}_{ij}$ share the same sparsity pattern across sample groups, the corresponding regression coefficients $\beta^{(\mathcal{S},l)}_{\cdot,i}$ also exhibit a common support across the $m$ sample groups, albeit with potentially distinct values. To harness this group sparsity assumption, we treat the coefficients $\beta^{(\mathcal{S},1)}_{ji},\dots,\beta^{(\mathcal{S},m)}_{ji}$ as a group of parameters for each pair $(i,j)$ and apply group Lasso (Yuan and Lin, 2006) to improve estimation by borrowing strength across sample groups:
[…]
@@
Once the regression coefficients are estimated, the residuals are computed as $\widehat{\epsilon}^{(\mathcal{S},k,l)}_{ti}:=X^{(k,l)}_{ti}-X^{(k,l)}_{t,\cdot}\widehat{\beta}^{(\mathcal{S},l)}_{\cdot,i}$. Although the empirical covariance matrix of the fitted residuals provides a straightforward estimate of $\Phi_{ij}^{(\mathcal{S},l)}$, this estimate is biased due to the lasso-type penalty, leading to a larger error rate than the expected $1/\sqrt{n_{l}p}$ rate (Liu, 2013). To address this bias, we adopt a bias-correction term in the covariance estimate:
@@
We then estimate the spatial precision matrix $\Omega^{(\mathcal{S},l)}$ and the partial correlation matrix $\rho^{(\mathcal{S},l)}$ as follows:
@@
While our partial correlation estimator is similar to the form proposed by Chen and Liu (2018) for a single matrix-variate Gaussian graphical model, our approach leverages information from multiple sample groups using the group Lasso estimate, leading to a faster convergence rate, as demonstrated in Theorem 4.1, which in turn facilitates the subsequent inference procedure.